OpenStudy (anonymous):

ASAP plz Just the answer A child is creating a pyramid with building blocks. The top three levels include 5 blocks, 12 blocks, and 19 blocks. Part 1: How many blocks would be needed for a pyramid 25 levels tall? (5 points) Part 2: Use complete sentences to explain how a sum of an arithmetic series was applied.

OpenStudy (anonymous):

@Jim766 plz help

OpenStudy (anonymous):

@kausarsalley help

OpenStudy (anonymous):

sequence and series

OpenStudy (anonymous):

good i got it without even reading the b

OpenStudy (anonymous):

medal to correct answer

OpenStudy (anonymous):

the terms are 5,12,19 and the common difference is 7

OpenStudy (anonymous):

can i have the formula for arithmetic progression

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

\[a_n=a_1+(n-1)d\]

OpenStudy (anonymous):

ok whats the answer plz

OpenStudy (anonymous):

plz asap

OpenStudy (anonymous):

need this

OpenStudy (anonymous):

OOK GIVE ME A MINUTE

OpenStudy (anonymous):

there you go buddy.

OpenStudy (anonymous):

no that different numbers

OpenStudy (anonymous):

come on i need this plz ASAP

OpenStudy (anonymous):

did you see the solution \[f(1)=5\]

OpenStudy (anonymous):

f(3)=12+7

OpenStudy (anonymous):

“You see, in this world there’s two kinds of people, my friend: Those with loaded guns, and those who dig. You dig.”

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

f(25)=19+(22*7)=

OpenStudy (anonymous):

i think that should give you the answer @pvs285

OpenStudy (anonymous):

i am not sure i just looked up what the other guy did

OpenStudy (anonymous):

dont worry @pvs285 i think @Luis_Rivera was talking to me

OpenStudy (anonymous):

i need that answer

OpenStudy (anonymous):

@satellite73 Help

OpenStudy (anonymous):

173 buddy may be you shuld attempt the algebra

OpenStudy (anonymous):

but part 2

OpenStudy (anonymous):

part 2 help

OpenStudy (anonymous):

@Jim766 help

OpenStudy (anonymous):

f(1)=5 f(n)=f(n-1)+d

OpenStudy (anonymous):

complete sentences

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

in maths formulas are sentences

OpenStudy (anonymous):

come on really

OpenStudy (anonymous):

@christianfilms11 come on

OpenStudy (anonymous):

I am sorry! that's all I got :'(

OpenStudy (anonymous):

jim help

OpenStudy (jim766):

Arithmetic seq An = a1 + (n-1)d formual for arth seq A25 = a1 + (25-1)(-7) 25 rows, comm dif = -7 5 = a1 + (24)(-7) 5 = a1 + -168 173 = a1 the bottom row in the pyrimid would have 173 They tell you that the top 3 rows have 5, 12, 19 That means top row , A25, has 5

OpenStudy (jim766):

Need to do this to find A1 Sum of arith seq is \[sn = \frac{ n(a1 + an) }{ 2 }\] n is the number of rows 25 A1 is 1st row 173 An is 25th row 5